OpenStudy (anonymous):

im (x+1)^7(2x)^2 x->2 i found the answer to this and its asking now what laws did i use? quotient, power, product, difference, root, constant mult, and sum?

OpenStudy (amistre64):

does x=2 make it go bad?

OpenStudy (amistre64):

as far as the laws go, i got no idea what you did, or even which ones apply :)

OpenStudy (anonymous):

no the lim is 17496 i just dont know the rules i used to get that lol

OpenStudy (amistre64):

im sure you added, multiploed and powered

OpenStudy (anonymous):

lol me too but its not taking that as an answer

myininaya (myininaya):

how did you get that

OpenStudy (anonymous):

i just plugged in 2 and got the answer lol

myininaya (myininaya):

\[\lim_{x \rightarrow 2}(x+1)^7(2x)^2=(2+1)^7(2(2))^2=(3)^74^2=2187(16)=34992\]

myininaya (myininaya):

thats what you get pluggin in 2

OpenStudy (anonymous):

Plug in the the limit that x is approaching and solve like normal

OpenStudy (anonymous):

oh ya your right sorry i typed it in wrong. it was supposto be 2x^2. so you would do 2187*8

myininaya (myininaya):

yes now you get what you got

myininaya (myininaya):

\[\lim_{x \rightarrow 2}[(x+1)^72x^2]=\lim_{x \rightarrow 2}(x+1)^7 \cdot \lim_{x \rightarrow 2}(2x^2)\] \[=(\lim_{x \rightarrow 2}(x+1))^7 \cdot 2 \lim_{x \rightarrow 2}x^2\] \[=(2+1)^7 \cdot 2 \cdot (2)^2\]

myininaya (myininaya):

product used in first step power and constant used in second step

myininaya (myininaya):

and i guess you could put another step inside the parentheses and used the sum limit rule

OpenStudy (anonymous):

why dont you just say its a continuous function defined over all x and thus the limit is equal to the function value

myininaya (myininaya):

he said he wanted limit properties to be used